# Local Asymptotic Normality for Mixed Fractional Brownian Motion under High-Frequency Observation

### v2

@@ 4.1 Projection and exact orthogonality
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma\mathbf{a}_{n}}{2}\,S_{\sigma,n}.$ (43)
@@
Write $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}(Z^{\top}C_{n}Z-\mathrm{tr}C_{n})$ and $R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}D_{n}^{\perp})$. By Wick’s identity for centered Gaussian quadratic forms,
@@ 4.2 Asymptotics of the projection coefficient and the orthogonal Frobenius norm
$\mathrm{tr}(C_{n}D_{n})=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})\,J_{0}(H,\sigma)+J_{1}(H,\sigma)+o\!\big(\ln(1/\Delta_{n})\big)\Bigr),$ (44)
@@
On the domain $\{\eta<|\lambda|\leq\pi\}$, the functions $f_{H}$ and $\dot{f}_{H}$ are bounded, and $a_{n}(\lambda)\geq 1$. Thus, $|g_{n}(\lambda)h_{n}(\lambda)|\leq C_{\eta}$ uniformly in $n$, implying $I_{1,n}^{(\infty)}=O(1)$. Since $p>1/2$, $\Delta_{n}^{1-2p}\to\infty$, so $I_{1,n}^{(\infty)}=o(\Delta_{n}^{1-2p}\ln(1/\Delta_{n}))$.
@@
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)=\Delta_{n}^{1-2p}w(x)\Bigl(2\ln(1/\Delta_{n})+C_{H}-2\ln|x|\Bigr)+o\!\big(\Delta_{n}^{1-2p}\ln(1/\Delta_{n})\big).$ (47)
By Lemma 4, the integrand is dominated by an integrable function involving logarithmic factors. Thus, we can apply the Dominated Convergence Theorem. Integrating the leading terms over $\mathbb{R}$ (since $\eta/\Delta_{n}\to\infty$) yields:
$I_{1,n}^{(0)}$ $=\Delta_{n}^{1-2p}\left[2\ln(1/\Delta_{n})\int_{\mathbb{R}}w(x)dx+\int_{\mathbb{R}}w(x)(C_{H}-2\ln|x|)dx+o(\ln(1/\Delta_{n}))\right]$
$=\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})J_{0}+J_{1}+o(\ln(1/\Delta_{n}))\Bigr).$
@@
Insert (33), (44), (37), and use $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+m+o(1)$ with $m=J_{1}/J_{0}$ from Lemma 11. A direct algebraic cancellation shows that the coefficients of $\ln^{2}(1/\Delta_{n})$ and $\ln(1/\Delta_{n})$ vanish, and the remaining constant term equals $J_{2}-m^{2}J_{0}=J_{2}-J_{1}^{2}/J_{0}$. This yields (49). Positivity follows since $J_{\perp}=\int_{\mathbb{R}}w(x)(q(x)-m)^{2}dx$ with $q(x)=C_{H}-2\ln|x|$ not $w$-a.e. constant. ∎
@@
From Lemma 11 we have $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+O(1)$, hence $|\mathbf{a}_{n}|=O(\log n)$ because $\Delta_{n}=n^{-\alpha}$. Together with Lemma 7 this yields
@@ 4.3 Joint CLT after multiplying by the rate matrix
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\qquad M_{n}^{(2)}:=\begin{pmatrix}1&0\\ -\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix},\qquad M_{n}:=M_{n}^{(2)}M_{n}^{(1)}=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})-\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix}$. (50)
@@
$R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})\Big).$
@@
$\frac{Z^{\top}D_{n}^{\perp}Z-\mathrm{tr}(D_{n}^{\perp})}{\sqrt{2}\left\lVert D_{n}^{\perp}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
@@
$u_{1}\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}+u_{2}\frac{R_{H,n}^{\perp}}{\sqrt{\mathcal{T}_{n}}}=\frac{\gamma_{n}}{\sqrt{\mathcal{T}_{n}}}\Big(Z^{\top}M_{n,u}Z-\mathrm{tr}(M_{n,u})\Big),\qquad M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n}^{\perp}.$
@@
and $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}\leq|u_{1}|\sigma^{-1}\left\lVert C_{n}\right\rVert_{\mathrm{op}}+|u_{2}|2^{-1}\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}$. Since $\left\lVert M_{n,u}\right\rVert_{F}^{2}$ is a nonnegative sum of the two terms above, we obtain $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert M_{n,u}\right\rVert_{F}\to 0$ from $\left\lVert C_{n}\right\rVert_{\mathrm{op}}/\left\lVert C_{n}\right\rVert_{F}\to 0$ (verified in Proposition 3) and Lemma 13. Applying Lemma 2 to $M_{n,u}$ yields a one-dimensional normal limit for every $u$, hence joint convergence by Cramér–Wold.
@@ 5.1 Matrix Taylor Expansions
To establish the Local Asymptotic Normality, we need to analyze the asymptotic behavior of the log-likelihood ratio. This ratio involves non-linear matrix functions of the covariance matrix, specifically the logarithm of the determinant and the matrix inverse. To handle these terms, we rely on second-order Taylor expansions with respect to the perturbation matrix. The following lemma provides the fundamental matrix inequalities required to expand these terms and strictly control the remainder terms using the operator norm and the Frobenius norm (trace of the square).
@@ 5.2 Final Proof of Theorem 2
Fix $h\in\mathbb{R}^{2}$ and set $\theta_{h}:=\theta+hr_{n}^{-1}$. Write $V:=V_{n}(\theta)$ and $V_{h}:=V_{n}(\theta_{h})$. Under $\mathbb{P}_{\theta}$, the whitened vector $Z:=V^{-1/2}X_{n}$ is $\mathcal{N}(0,I_{n})$. Define the symmetric perturbation matrix
@@
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)=-\frac{1}{2}\log\det(I_{n}+S_{n}(h))-\frac{1}{2}\,Z^{\top}\Big((I_{n}+S_{n}(h))^{-1}-I_{n}\Big)Z.$ (51)
@@
$A_{n}(\theta_{h})-A_{n}(\theta)=(\gamma_{h}-\gamma)T_{n}(H)+\gamma(H_{h}-H)\dot{T}_{n}(H)+\mathcal{R}_{A,n},$
where $\mathcal{R}_{A,n}$ collects the second-order Taylor terms. Moreover, this remainder is negligible after sandwiching: using $\|\ddot{T}_{n}(H)\|_{\mathrm{op}}=O(\log^{2}n)$ (coming from the $\log^{2}|2\sin(\lambda/2)|$ singularity in $\partial_{H}^{2}\log f_{H}$) and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, we obtain $\|A_{n}(\theta)^{-1/2}\mathcal{R}_{A,n}A_{n}(\theta)^{-1/2}\|_{\mathrm{op}}=o(1)$ (since $|\log\Delta_{n}|\asymp\log n$ and $\mathcal{T}_{n}\to\infty$). Sandwiching by $A_{n}(\theta)^{-1/2}$ gives
@@
and and $\|R_{S,n}\|_{\mathrm{op}}=o(1)$ uniformly for fixed $h$. Using $\left\lVert C_{n}(\theta)\right\rVert_{\mathrm{op}}\leq 1/\gamma$ and $\left\lVert D_{n}(\theta)\right\rVert_{\mathrm{op}}=O(\gamma^{-1}\log n)$ (by Lemma 7), together with $(\gamma_{h}-\gamma)=O(\gamma\mathcal{T}_{n}^{-1/2})$ and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, we obtain
@@
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)$ $=\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)-\frac{1}{4}\mathrm{tr}\!\big(S_{n}(h)^{2}\big)$
$\quad-\frac{1}{2}\big(Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})\big)-\frac{1}{2}R_{\log}(S_{n}(h))-\frac{1}{2}Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z.$ (52)
@@
$\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)=\sum_{i=1}^{2}\delta_{n,i}\,\partial_{\theta_{i}}\ell_{n}(\theta)+\frac{1}{2}\big(Z^{\top}R_{n}Z-\mathrm{tr}R_{n}\big).$
The last term is $o_{\mathbb{P}}(1)$ since $\mathrm{Var}(Z^{\top}R_{n}Z-\mathrm{tr}R_{n})=2\mathrm{tr}(R_{n}^{2})=o(1)$. Hence, using $\delta_{n}^{\top}\nabla\ell_{n}(\theta)=h^{\top}(r_{n}^{-\top}\nabla\ell_{n}(\theta))$, we have
$\frac{1}{2}\big(Z^{\top}S_{n}(h)Z-\mathrm{tr}S_{n}(h)\big)=\delta_{n}^{\top}\nabla\ell_{n}(\theta)+o_{\mathbb{P}}(1)=h^{\top}\Xi_{n}+o_{\mathbb{P}}(1),$ (53)
@@
Finally, we control the remaining high-order error terms in (52) and show they converge to zero in probability. Using $\mathrm{Var}(Z^{\top}AZ-\mathrm{tr}A)=2\mathrm{tr}(A^{2})$ and $\mathrm{tr}(S^{4})\leq\left\lVert S\right\rVert_{\mathrm{op}}^{2}\mathrm{tr}(S^{2})$, we obtain
$\mathrm{Var}\big(Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})\big)=2\mathrm{tr}(S_{n}(h)^{4})\leq 2\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{2}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
by the results established above, hence $Z^{\top}S_{n}(h)^{2}Z-\mathrm{tr}(S_{n}(h)^{2})=o_{\mathbb{P}}(1)$. Moreover, Lemma 14 and (54) give
@@
Also, write $Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z=\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))+\big(Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)$. Since $R_{\mathrm{inv}}(S)=\sum_{k\geq 3}(-1)^{k}S^{k}$, we have $|\mathrm{tr}(R_{\mathrm{inv}}(S))|\leq\sum_{k\geq 3}|\mathrm{tr}(S^{k})|\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}\mathrm{tr}(S^{2})$, hence $|\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))|=o(1)$. Furthermore,
$\mathrm{Var}\big(Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)=2\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h))^{2})\leq C\,\mathrm{tr}(S_{n}(h)^{6})\leq C\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{4}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
so $Z^{\top}R_{\mathrm{inv}}(S_{n}(h))Z=o_{\mathbb{P}}(1)$.

### v3

@@ 4.1 Projection and exact orthogonality
$R_{H,n}^{\perp}:=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}^{\perp}Z_{n}-\mathrm{tr}(D_{n}^{\perp})\Big)=R_{H,n}-\frac{\sigma\mathbf{a}_{n}}{2}\,S_{\sigma,n}.$ (43)
@@
Write $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}(Z_{n}^{\top}C_{n}Z_{n}-\mathrm{tr}(C_{n}))$ and $R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}(Z_{n}^{\top}D_{n}^{\perp}Z_{n}-\mathrm{tr}D_{n}^{\perp})$. By Wick’s identity for centered Gaussian quadratic forms,
@@ 4.2 Asymptotics of the projection coefficient and the orthogonal Frobenius norm
$\mathrm{tr}(C_{n}D_{n})=\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})\,J_{0}(H,\sigma)+J_{1}(H,\sigma)+o(1)\Bigr),$ (44)
@@
On the domain $\{\eta<|\lambda|\leq\pi\}$, the functions $f_{H}$ and $\dot{f}_{H}$ are bounded, and $a_{n}(\lambda)\geq 1$. Thus, $|g_{n}(\lambda)h_{n}(\lambda)|\leq C_{\eta}$ uniformly in $n$, implying $I_{1,n}^{(\infty)}=O(1)$. Since $p>1/2$, $\Delta_{n}^{1-2p}\to\infty$, so $I_{1,n}^{(\infty)}=o(\Delta_{n}^{1-2p})$.
@@
Moreover, by the $C^{2}$-regularity of the low-frequency factor (Lemma 3), the convergence in (46) is uniform on compact $x$-sets and yields an error of smaller order than $\Delta_{n}^{1-2p}$ after integration (in particular, $\Delta_{n}^{2}\ln(1/\Delta_{n})\to 0$). More precisely, write $\kappa_{n}(x)=\Delta_{n}^{-2p}w(x)+r_{n}(x)$ on $|x|\leq\eta/\Delta_{n}$. Then, by the $C^{2}$-regularity in Lemma 3 and the same domination used in Lemma 6,
$\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}|r_{n}(x)|\,dx=o(\Delta_{n}^{1-2p}),\qquad\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}|r_{n}(x)|\,\ln(1/\Delta_{n})\,dx=o(\Delta_{n}^{1-2p}),$
since $\Delta_{n}^{2}\ln(1/\Delta_{n})\to 0$.
@@
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)=\Delta_{n}^{1-2p}w(x)\Bigl(2\ln(1/\Delta_{n})+C_{H}-2\ln|x|\Bigr)+o(\Delta_{n}^{1-2p}).$ (47)
By Lemma 4, the integrand is dominated by an integrable function up to a constant factor. Thus, we can apply the Dominated Convergence Theorem. Integrating the leading terms over $\mathbb{R}$ (since $\eta/\Delta_{n}\to\infty$) yields:
$I_{1,n}^{(0)}$ $=\Delta_{n}^{1-2p}\left[2\ln(1/\Delta_{n})\int_{\mathbb{R}}w(x)\,dx+\int_{\mathbb{R}}w(x)(C_{H}-2\ln|x|)dx+o(1)\right]$
$=\Delta_{n}^{1-2p}\Bigl(2\ln(1/\Delta_{n})J_{0}+J_{1}+o(1)\Bigr).$
@@
Using the definition $\mathbf{a}_{n}=\mathrm{tr}(C_{n}D_{n})/\mathrm{tr}(C_{n}^{2})$ in (42), this can be rewritten as
$\mathrm{tr}\big((D_{n}^{\perp})^{2}\big)=\mathrm{tr}(D_{n}^{2})-\frac{\mathrm{tr}(C_{n}D_{n})^{2}}{\mathrm{tr}(C_{n}^{2})}.$ (50)
By the trace approximation bounds in Lemma 5, the remainders in (21)–(23) are $o\!\big(n\Delta_{n}^{1-2p}\big)$, hence it suffices to analyze the corresponding integrals:
$I_{0,n}:=\int_{-\pi}^{\pi}g_{n}(\lambda)^{2}\,d\lambda,\qquad I_{1,n}:=\int_{-\pi}^{\pi}g_{n}(\lambda)h_{n}(\lambda)\,d\lambda,\qquad I_{2,n}:=\int_{-\pi}^{\pi}h_{n}(\lambda)^{2}\,d\lambda,$
where $g_{n}(\lambda)=f_{H}(\lambda)/(1+\gamma_{n}f_{H}(\lambda))$ and $h_{n}(\lambda)=\dot{f}_{H}(\lambda)/(1+\gamma_{n}f_{H}(\lambda))$. Indeed,
$\mathrm{tr}(D_{n}^{2})=\frac{n}{2\pi}I_{2,n}+o\!\big(n\Delta_{n}^{1-2p}\big),\qquad\mathrm{tr}(C_{n}D_{n})=\frac{n}{2\pi}I_{1,n}+o\!\big(n\Delta_{n}^{1-2p}\big),\qquad$
$\mathrm{tr}(C_{n}^{2})=\frac{n}{2\pi}I_{0,n}+o\!\big(n\Delta_{n}^{1-2p}\big).$
Consequently, (50) implies
$\mathrm{tr}\big((D_{n}^{\perp})^{2}\big)=\frac{n}{2\pi}\left(I_{2,n}-\frac{I_{1,n}^{2}}{I_{0,n}}\right)+o\!\big(n\Delta_{n}^{1-2p}\big).$
Fix $\eta\in(0,\pi)$ and split each $I_{k,n}=I_{k,n}^{(0)}+I_{k,n}^{(\infty)}$ over $|\lambda|\leq\eta$ and $\eta<|\lambda|\leq\pi$. On $\{\eta<|\lambda|\leq\pi\}$, $f_{H}$ and $\dot{f}_{H}$ are bounded and $a_{n}(\lambda)\geq 1$, so $I_{k,n}^{(\infty)}=O(1)$ for $k=0,1,2$, therefore these contributions are $O(1)$ and thus negligible compared with the leading order $\Delta_{n}^{1-2p}$. Moreover, although $I_{1,n}$ enters through the ratio $I_{1,n}^{2}/I_{0,n}$, the high-frequency parts still remain negligible: since $I_{0,n}^{(0)}\asymp\Delta_{n}^{1-2p}$ and $I_{1,n}^{(0)}=O(\Delta_{n}^{1-2p}\ln(1/\Delta_{n}))$, a perturbation $I_{k,n}=I_{k,n}^{(0)}+O(1)$ changes $I_{1,n}^{2}/I_{0,n}$ by at most $O(\ln(1/\Delta_{n}))=o(\Delta_{n}^{1-2p})$. Hence $I_{2,n}-I_{1,n}^{2}/I_{0,n}$ can be replaced by its low-frequency counterpart up to $o(\Delta_{n}^{1-2p})$.
On $|\lambda|\leq\eta$, set $\lambda=\Delta_{n}x$ and use the notation of Lemma 10:
Then
$I_{0,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\kappa_{n}(x)\,dx,\quad I_{1,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\kappa_{n}(x)\,b_{n}(x)\,dx,$
$I_{2,n}^{(0)}=\Delta_{n}\int_{|x|\leq\eta/\Delta_{n}}\kappa_{n}(x)\,b_{n}(x)^{2}\,dx.$
By Lemma 11, $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+m+o(1)$ with $m=J_{1}/J_{0}$, while $b_{n}(x)=2\ln(1/\Delta_{n})+q(x)+o(1)$ pointwise for fixed $x\neq 0$, where $q(x)=C_{H}-2\ln|x|$. Moreover, $\kappa_{n}(x)=\Delta_{n}^{-2p}w(x)(1+o(1))$ pointwise with $w$ given in (27), and the integrands are dominated as in Lemma 4. Therefore, by dominated convergence and $\eta/\Delta_{n}\to\infty$,
$I_{2,n}^{(0)}-\frac{(I_{1,n}^{(0)})^{2}}{I_{0,n}^{(0)}}=\Delta_{n}^{1-2p}\left(\int_{\mathbb{R}}w(x)\,(q(x)-m)^{2}\,dx+o(1)\right)=\Delta_{n}^{1-2p}\big(J_{\perp}+o(1)\big).$
Combining the low- and high-frequency parts yields (49). Positivity follows since $q$ is not $w$-a.e. constant. ∎
@@
From Lemma 11 we have $\mathbf{a}_{n}=2\ln(1/\Delta_{n})+O(1)+o(1)$, hence $|\mathbf{a}_{n}|=O(\log n)$ because $\Delta_{n}=n^{-\alpha}$. Together with Lemma 7 this yields
@@ 4.3 Joint CLT after multiplying by the rate matrix
$M_{n}^{(1)}:=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})&1\end{pmatrix},\quad M_{n}^{(2)}:=\begin{pmatrix}1&0\\ -\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix},\quad M_{n}:=M_{n}^{(2)}M_{n}^{(1)}=\begin{pmatrix}1&0\\ -\sigma\ln(\Delta_{n})-\frac{\sigma\mathbf{a}_{n}}{2}&1\end{pmatrix}$. (51)
@@
$R_{H,n}^{\perp}=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}^{\perp}Z_{n}-\mathrm{tr}(D_{n}^{\perp})\Big).$
@@
$\frac{Z_{n}^{\top}D_{n}^{\perp}Z_{n}-\mathrm{tr}(D_{n}^{\perp})}{\sqrt{2}\left\lVert D_{n}^{\perp}\right\rVert_{F}}\Rightarrow\mathcal{N}(0,1).$
@@
$u_{1}\frac{S_{\sigma,n}}{\sqrt{\mathcal{T}_{n}}}+u_{2}\frac{R_{H,n}^{\perp}}{\sqrt{\mathcal{T}_{n}}}=\frac{\gamma_{n}}{\sqrt{\mathcal{T}_{n}}}\Big(Z_{n}^{\top}M_{n,u}Z_{n}-\mathrm{tr}(M_{n,u})\Big),\qquad M_{n,u}:=u_{1}\sigma^{-1}C_{n}+u_{2}\tfrac{1}{2}D_{n}^{\perp}.$
@@
and $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}\leq\sigma^{-1}|u_{1}|\left\lVert C_{n}\right\rVert_{\mathrm{op}}+2^{-1}|u_{2}|\left\lVert D_{n}^{\perp}\right\rVert_{\mathrm{op}}$. Since $\left\lVert M_{n,u}\right\rVert_{F}^{2}$ is a nonnegative sum of the two terms above, we obtain $\left\lVert M_{n,u}\right\rVert_{\mathrm{op}}/\left\lVert M_{n,u}\right\rVert_{F}\to 0$ from $\left\lVert C_{n}\right\rVert_{\mathrm{op}}/\left\lVert C_{n}\right\rVert_{F}\to 0$ (verified in Proposition 3) and Lemma 13. Applying Lemma 2 to $M_{n,u}$ yields a one-dimensional normal limit for every $u$, hence joint convergence by Cramér–Wold.
@@
Remark 3.
In the matrix $M_{n}$, we can change $\mathbf{a_{n}}$ directly with $2\ln\!\Big(\frac{1}{\Delta_{n}}\Big)+\frac{J_{1}}{J_{0}}$ and obtain the same result.
@@ 5.1 Matrix Taylor Expansions
To establish the Local Asymptotic Normality, we need to analyze the asymptotic behavior of the log-likelihood ratio. This ratio involves non-linear matrix functions of the covariance matrix, specifically the logarithm of the determinant and the matrix inverse. To handle these terms, we rely on second-order Taylor expansions with respect to the perturbation matrix. The following lemma provides the fundamental matrix inequalities required to expand these terms and strictly control the remainder terms using the operator norm and the Frobenius norm.
@@ 5.2 Final Proof of Theorem 2
Fix $h\in\mathbb{R}^{2}$ and set $\theta_{h}:=(\sigma_{h},H_{h})=\theta+hr_{n}^{-1}$. Write $V:=V_{n}(\theta)$ and $V_{h}:=V_{n}(\theta_{h})$. Under $\mathbb{P}_{\theta}$, the whitened vector $Z_{n}:=V^{-1/2}X_{n}$ is $\mathcal{N}(0,I_{n})$. Define the symmetric perturbation matrix
@@
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)=-\frac{1}{2}\log\det(I_{n}+S_{n}(h))-\frac{1}{2}\,Z_{n}^{\top}\Big((I_{n}+S_{n}(h))^{-1}-I_{n}\Big)Z_{n}.$ (52)
@@
$A_{n}(\theta_{h})-A_{n}(\theta)=(\gamma_{h}-\gamma)T_{n}(H)+\gamma(H_{h}-H)\dot{T}_{n}(H)+R_{A,n},$
where $\gamma_{h}:=\gamma(\theta_{h})$ and $R_{A,n}$ collects the second-order Taylor terms. Moreover, this remainder is negligible after sandwiching: using $\|\ddot{T}_{n}(H)\|_{\mathrm{op}}=O(\log^{2}n)$ (coming from the $\log^{2}|2\sin(\lambda/2)|$ singularity in $\partial_{H}^{2}\log f_{H}$) and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, since $|\log\Delta_{n}|\asymp\log n$ and $\mathcal{T}_{n}\to\infty$), we obtain
$\|A_{n}(\theta)^{-1/2}R_{A,n}A_{n}(\theta)^{-1/2}\|_{\mathrm{op}}=o(1).$
Sandwiching by $A_{n}(\theta)^{-1/2}$ gives
@@
and $\|R_{S,n}\|_{\mathrm{op}}=o(1)$ uniformly for fixed $h$. Using $\left\lVert C_{n}(\theta)\right\rVert_{\mathrm{op}}\leq 1/\gamma$ and $\left\lVert D_{n}(\theta)\right\rVert_{\mathrm{op}}=O(\gamma^{-1}\log n)$ (by Lemma 7), together with $(\gamma_{h}-\gamma)=O(\gamma\mathcal{T}_{n}^{-1/2})$ and $H_{h}-H=O\big(\mathcal{T}_{n}^{-1/2}|\log\Delta_{n}|^{-1}\big)$, we obtain
@@
$\ell_{n}(\theta_{h})-\ell_{n}(\theta)$ $=\frac{1}{2}\big(Z_{n}^{\top}S_{n}(h)Z_{n}-\mathrm{tr}S_{n}(h)\big)-\frac{1}{4}\mathrm{tr}\!\big(S_{n}(h)^{2}\big)$
$\quad-\frac{1}{2}\big(Z_{n}^{\top}S_{n}(h)^{2}Z_{n}-\mathrm{tr}(S_{n}(h)^{2})\big)-\frac{1}{2}R_{\log}(S_{n}(h))-\frac{1}{2}Z_{n}^{\top}R_{\mathrm{inv}}(S_{n}(h))Z_{n}.$ (53)
@@
$\frac{1}{2}\big(Z_{n}^{\top}S_{n}(h)Z_{n}-\mathrm{tr}S_{n}(h)\big)=\sum_{i=1}^{2}\delta_{n,i}\,\partial_{\theta_{i}}\ell_{n}(\theta)+\frac{1}{2}\big(Z_{n}^{\top}R_{n}Z_{n}-\mathrm{tr}(R_{n})\big).$
The last term is $o_{\mathbb{P}}(1)$ since $\mathrm{Var}(Z_{n}^{\top}R_{n}Z_{n}-\mathrm{tr}(R_{n}))=2\mathrm{tr}(R_{n}^{2})=o(1)$. Hence, using $\delta_{n}^{\top}\nabla\ell_{n}(\theta)=h^{\top}(r_{n}^{-\top}\nabla\ell_{n}(\theta))$, we have
$\frac{1}{2}\big(Z_{n}^{\top}S_{n}(h)Z_{n}-\mathrm{tr}S_{n}(h)\big)=\delta_{n}^{\top}\nabla\ell_{n}(\theta)+o_{\mathbb{P}}(1)=h^{\top}\Xi_{n}+o_{\mathbb{P}}(1),$ (54)
@@
Finally, we control the remaining high-order error terms in (53) and show they converge to zero in probability. Using $\mathrm{Var}(Z_{n}^{\top}AZ_{n}-\mathrm{tr}(A))=2\mathrm{tr}(A^{2})$ and $\mathrm{tr}(S^{4})\leq\left\lVert S\right\rVert_{\mathrm{op}}^{2}\mathrm{tr}(S^{2})$, we obtain
$\mathrm{Var}\big(Z_{n}^{\top}S_{n}(h)^{2}Z_{n}-\mathrm{tr}(S_{n}(h)^{2})\big)=2\mathrm{tr}(S_{n}(h)^{4})\leq 2\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{2}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
by the results established above, hence $Z_{n}^{\top}S_{n}(h)^{2}Z_{n}-\mathrm{tr}(S_{n}(h)^{2})=o_{\mathbb{P}}(1)$. Moreover, Lemma 14 and (55) give
@@
Also, write $Z_{n}^{\top}R_{\mathrm{inv}}(S_{n}(h))Z_{n}=\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))+\big(Z_{n}^{\top}R_{\mathrm{inv}}(S_{n}(h))Z_{n}-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)$. Since $R_{\mathrm{inv}}(S)=\sum_{k\geq 3}(-1)^{k}S^{k}$, we have $|\mathrm{tr}(R_{\mathrm{inv}}(S))|\leq\sum_{k\geq 3}|\mathrm{tr}(S^{k})|\leq C\,\left\lVert S\right\rVert_{\mathrm{op}}\mathrm{tr}(S^{2})$, hence $|\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))|=o(1)$. Furthermore,
$\mathrm{Var}\big(Z_{n}^{\top}R_{\mathrm{inv}}(S_{n}(h))Z_{n}-\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h)))\big)$ $=2\mathrm{tr}(R_{\mathrm{inv}}(S_{n}(h))^{2})$
$\leq C\,\mathrm{tr}(S_{n}(h)^{6})\leq C\,\left\lVert S_{n}(h)\right\rVert_{\mathrm{op}}^{4}\,\mathrm{tr}(S_{n}(h)^{2})\to 0,$
so $Z_{n}^{\top}R_{\mathrm{inv}}(S_{n}(h))Z_{n}=o_{\mathbb{P}}(1)$.